find the slope of any line perpendicular to the line through points(0,5) and
(-3,-4) I started with y=0(x)+ 5, y=-3x-4 i am not sure what to do next.

Answer by venugopalramana(3286)

find the slope of any line perpendicular to the line through points(0,5) and
(-3,-4) I started with y=0(x)+ 5, y=-3x-4 i am not sure what to do next.
THAT IS NOT THE WAY FIRST FIND SLOPE OF THE LINE JOINING THE 2 POINTS ..IT IS..(Y2-Y1)/(X2-X1)=(5+4)/(0+3)=9/3=3
NOW SLOPE OF THE PERPENDICULAR LINE = -1/3
RULE FOR PERPENDICULAR LINES =PRODUCT OF SLOPES = -1